OpenStudy (anonymous):

PB is a line segment on a number line. It has endpoints at -2 and 12. What is the coordinate of its midpoint.

OpenStudy (firekat97):

so we know that a number line just exists on the horizontal plane, so we know that our two points are x = =2 and x = 12, so the mid point can actually just be found by averaging these two points, so average = ((-2) + (12))/2 so = (10)/2 = 5 therefore the midpoint is at 5

OpenStudy (anonymous):

thanks!